OpenStudy (anonymous):

solve 2(2-2x)=-3x A.-4 B.4 C.6 D.0

OpenStudy (anonymous):

well so \( 4 - 4x = -3x \)

OpenStudy (anonymous):

and if u add \( 4x\) to the both sides u'll get \( 4=4x - 3x \)

OpenStudy (anonymous):

got it? can u find the answer now?

OpenStudy (anonymous):

\[4-4x=-3x\]\[-4x+3x=-4\]\[x=?\] solve it @BleuSoul57

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

\[-x=-4\]\[x=\frac{ -4 }{ -1 }\]\[x=?\]

OpenStudy (anonymous):

solve it @BleuSoul57

OpenStudy (anonymous):

hint,\[-1\times-1=1\] \[\frac{ -1 }{ -1 }=1\] @BleuSoul

OpenStudy (anonymous):

But there's no 1

OpenStudy (anonymous):

It's an example @BleuSoul57

OpenStudy (anonymous):

IT'S 4!

OpenStudy (anonymous):

Yes ! @BleuSoul57